OpenStudy (anonymous):

Xavier exercises by riding his bike. During his first week of training, he rides a total of 15 miles. Each week he increases the number of miles biked by 0.4 miles. How many total weeks will it take Xavier to first bike more than 500 miles total? a.8 b.15 c.19 d.26

OpenStudy (anonymous):

Each week he rides a total of 15 + .04(Number of weeks, I'll just use "n" for now on) Therefore, 500= S(15+0.4(n-1)), where s is sigma (that sideways m thingy), starting at x=1. You're solving for the number on top of sigma, so to put it in a less complex way: 500=15n + .04(n-1) (another way of phrasing this, I can't really explain the math without a proper math keyboard, but essentially it follows the definition of a series) This yeilds I believe 32.49 weeks, but that's illogical for this problem, so you round up so that the total number of whole weeks it takes for Xavier to bike at least 500 miles is 33 weeks. Hope this helps

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (anonymous):

ya thats really annoying can u guys please help me i don't get this and i need to turn it in soon

OpenStudy (boldjon):

is it a test?

OpenStudy (boldjon):

/quiz

OpenStudy (anonymous):

quiz

OpenStudy (alivejeremy):

hey need help

OpenStudy (anonymous):

\[S _{n}=\frac{ n }{ 2 }\left\{ 2a+\left( n-1 \right)d \right\}\] a=15 d=0.4 \[S _{n}=500\] find n

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

make a quadratic and then find n

OpenStudy (mom.):

mom is here ;)

OpenStudy (mom.):

who wants help ?<33

OpenStudy (anonymous):

\[1000=n \left\{ 2*15+\left( n-1 \right)\frac{ 4 }{ 10 } \right\}=n \left\{ \frac{ 150+2n-2 }{ 5 } \right\}\] 5000=n[148+2n] \[2n^2+148n-5000=0\] \[n^2+74n-2500=0\] \[n=\frac{ -74\pm \sqrt{74^2-4*1*-2500} }{ 2*1 }=\frac{ -74\pm 2 \sqrt{37^2+2500} }{ 2 }\] \[=-37\pm \sqrt{1369+2500}=-37\pm \sqrt{3869}=-37\pm62.2=25.2\] taking only positive sign n=26 (round)